# RefDrone: A Challenging Benchmark for Referring Expression Comprehension in Drone Scenes

### v2

@@ 5.1 Experimental results
Table 5: Experimental results on RSVG [56] dataset. Asterisk (*) denotes results reported in the EarthGPT [58] paper.
@@
Methods Pr@0.5 $\uparrow$ Pr@0.6 $\uparrow$ Pr@0.7 $\uparrow$ Pr@0.8 $\uparrow$ TransVG* [11] 72.41 67.38 60.05 49.10 MGVLF* [56] 76.78 72.68 66.74 56.42 EarthGPT* [58] 76.65 71.93 66.52 56.53 GDINO-T [34] 76.99 75.81 73.36 66.77 NGDINO-T 77.16 76.03 73.81 67.84
@@
Number Prediction Head Number Cross-Attention F1inst. Accinst. F1img. Accimg. FPS 30.50 19.02 29.65 21.07 13.5 30.62 19.09 29.90 21.18 12.8 31.25 19.51 30.73 21.71 12.9 33.34 20.98 32.45 22.84 12.3
Length F1inst. Accinst. F1img. Accimg. Params 1 31.63 19.80 31.23 22.12 1.58M 10 33.34 20.98 32.45 22.84 1.65M 100 32.06 20.10 32.23 22.81 2.34M
Methods F1inst. Accinst. F1img. Accimg. ReCLIP [44] 24.62 14.04 11.58 6.15 GPT4-o 52.38 35.65 35.50 22.38 RDAnnotator 58.14 41.13 37.07 23.54
5.2 Ablation studies
Ablation study on NGDINO components. Table 6 presents the analysis of each component in NGDINO. The number prediction head alone shows a slight increase in Accinst. from 19.02% to 19.09%, indicating the minimal impact of this auxiliary head on REC tasks. When introducing the number cross-attention without number selection, we observe a more substantial improvement, with Accinst. increasing to 19.51%. This improvement can be partially attributed to the additional parameters introduced in the decoder. The most significant improvement is achieved through the combination of both the number prediction task and the number cross-attention components. This contribution increases Accinst. to 20.98%. While the additional computational cost results in a minor decrease in inference speed (FPS), this trade-off yields a 1.96% increase in Accinst..
Ablation study on query length. Table 8 analyzes the impact of varying the selected number query length. Utilizing a minimal query length of 1 lacks the capacity to capture complex numerical information. Conversely, extending the query length to 100 increases parameter count and computational overhead, potentially leading to optimization challenges. Through these experiments, we determine that a query length of 10 provides an optimal trade-off.
Performance on number prediction. The number prediction head achieves an overall number prediction accuracy of 53.5%. During the bounding box prediction stage, this accuracy is observed to be 22.84%. This performance gap demonstrates the effectiveness of the number prediction head. Besides, the number prediction achieves a mean absolute error (MAE) of 0.51, suggesting high precision as predictions closely align with ground truth values.
Performance of two-stage instance ranking method. Table 8 analyzes our annotation framework RDAnnotator (without human verification) against alternative two-stage instance ranking methods for REC tasks. We compare: (1) GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ using only Step 3 of RDAnnotator, and (2) ReCLIP [44], which employs CLIP [40] for instance ranking. All experiments utilize Faster-RCNN [41] as the object detector (18.0 mAP on VisDrone2019-DET [60]). Results show RDAnnotator’s ability to produce high-quality annotations. Furthermore, superior instance-level metrics over end-to-end methods reveal the promising potential of two-stage ranking approaches in multi-target challenges.
@@ 5.3 Limitations
While NGDINO addresses the multi-target and no-target challenges, several challenges remain. Figure 7 shows some typical NGDINO failure cases, highlighting challenges in RefDrone. The examples show rich contextual reasoning, challenging backgrounds, and small-scale object detection. These challenging cases reflect real applications and highlight areas for future improvement.
@@ 6 Conclusion
In this work, we introduce RefDrone, a challenging benchmark specifically designed for referring expression comprehension in drone scenes. The dataset is constructed using the RDAnnotator framework, an efficient semi-automated annotation pipeline combining LMM-based modules and human oversight. RDAnnotator provides high-quality annotations while reducing human effort. Furthermore, we develop NGDINO to address the multi-target and no-target challenges in RefDrone. Experimental results demonstrate consistently low performance across existing approaches, underscoring the inherent difficulty of the RefDrone benchmark. In the future, we aim to further enhance NGDINO to address additional challenges presented by RefDrone. We also plan to extend RefDrone to larger-scale scenarios and introduce additional tasks such as referring expression segmentation and referring expression tracking. We believe RefDrone will serve as a valuable benchmark for advancing research in drone-based REC tasks.
@@ Appendix A Appendix
  1. 1.
    Details of the baseline method. (Appendix A.1)
  2. 2.
@@
  3. 3.
    Color attribute analysis of RefDrone. (Appendix A.3)
  4. 4.
    Additional dataset examples. (Appendix A.4)
  5. 5.
    Results with respect to different object scales. (Appendix A.5)
  6. 6.
    Results on RefCOCO/+/g datasets. (Appendix A.6)
  7. 7.
    Visualization results analysis. (Appendix A.7)
  8. 8.
    Prompts and examples in RDAnnotator. (Appendix A.8)
@@ A.1 Details of baseline methods
  • Large multimodal models:
    • Kosmos-2: 1.6B parameters.
    • ONE-PEACE: 4B parameters, visual grounding API.
    • Shikra: 7B parameters, delta-v1 version.
    • MiniGPT-v2: 7B parameters.
    • LLaVA: 7B parameters, v1.5 version.
    • Qwen-VL: 7B parameters.
    • Ferret: 7B parameters.
    • CogVLM: 7B parameters, grounding-specific version.
    • SPHINX-v2: 13B parameters.
    • Griffon: 13B parameters.
@@ A.2 Implementation details
NGDINO implementation. The NGDINO adopts a two-stage procedure to maintain training stability. First, we pre-train the number prediction head on the RefDrone dataset while initializing other components with GDINO [34] parameters. We then fine-tune the entire model. We also evaluate it on the gRefCOCO [28] dataset, which includes no-target and limited multi-target samples. Additionally, we evaluate it on RSVG [56], an aerial-view dataset with multi-target challenges.
Zero-shot evaluation details. For zero-shot evaluation, we use the original model checkpoints as provided in their respective papers. The implementations and checkpoints for GLIP and GDINO are obtained from MMDetection [6].
Fine-tuning evaluation details. Our fine-tuning protocol maintains consistency across all experiments to ensure a fair comparison. We preserve the original learning strategies while excluding random crop augmentation due to its negative effect on position-sensitive samples. For LMMs, we employ the LoRA [17] fine-tuning strategy and follow the instruction tuning data structures. All fine-tuning experiments are run for 5 epochs using 4 NVIDIA A100 GPUs.
@@
We design a hybrid color extraction pipeline consisting of two main components: (1) a WideResNet-101 classifier [55] trained on HSV color space-based labels. The color labels are refined via manual validation, ensuring robustness to annotation noise and label ambiguity. (2) An LMM-based color consistency module that validates color predictions through structured prompts. The latter cross-checks the classifier’s outputs to mitigate ambiguities caused by illumination variance or partial occlusion.
In the RefDrone dataset, color attributes are present in 69% of expressions. We initially use the HSV color space to define six primary colors: white, black, red, blue, green, and yellow. During the annotation process, RDAnnotator expands the color set by adding six more colors: orange, pink, grey, purple, brown, and silver. The distribution of these color terms across the expressions is illustrated in Figure 8.
@@
A.4 Dataset examples
@@
A.5 The results of different object scales.
We add small/medium/large instance accuracy metrics (ACCs, ACCm, ACCl) for the RefDrone dataset in Table 9.
Table 9: Fine-tuned results with respect to different object scales.
Methods Accs Accm Accl Methods Accs Accm Accl
GLIP 2.65 18.28 27.29 LLaVA 4.18 7.26 11.99
GDINO-T 8.56 20.66 33.99 Qwen-VL 0.73 10.60 30.18
NGDINO-T 10.84 23.38 40.13 RDAnnotator 32.54 47.97 47.17
A.6 Results on RefCOCO/+/g datasets.
Since the RefCOCO, RefCOCO+, RefCOCOg datasets contain only one instance per expression, the proposed NGDINO leverages the number branch primarily to address multi-instance and no-instance scenarios. As a result, the performance of NGDINO is relatively similar to that of GDINO on these datasets.
Table 10: Results on RefCOCO/+/g datasets.
RefCOCO RefCOCO+ RefCOCOg
TestA TestB TestA TestB Val Test
MDETR 90.4 82.67 85.52 72.96 83.35 83.31
GDINO-T 91.4 86.6 87.5 74.0 85.5 85.8
NGDINO-T 91.5 86.5 87.8 74.7 85.3 85.8
A.7 Visualization results compared NGDINO with baseline GDINO.
Figure 9 shows visualization results comparing NGDINO and the baseline GDINO on the proposed RefDrone dataset, demonstrating NGDINO’s effectiveness with multi-target samples.
Figure 9: Visualization results comparing (a) GDINO [34] and (b) NGDINO, where red boxes indicate true positives and yellow boxes denote false negatives.
A.8 Prompts and examples for RDAnnotator

### v3

@@ 5.3 Ablation studies
Table 5: Experimental results on RSVG [zhan2023rsvg] dataset. Asterisk (*) denotes results reported in the EarthGPT paper. Pr@0.5: percentage of predictions at IoU $\geq$ 0.5..
@@
Methods Pr@0.5 Pr@0.6 Pr@0.7 Pr@0.8 Pr@0.9 TransVG* [deng2021transvg] 72.41 67.38 60.05 49.10 27.84 MGVLF* [zhan2023rsvg] 76.78 72.68 66.74 56.42 35.07 EarthGPT* [zhang2024earthgpt] 76.65 71.93 66.52 56.53 37.63 GDINO-T [gdino] 76.99 75.81 73.36 66.77 49.96 NGDINO-T 77.16 76.03 73.81 67.84 51.39
@@
Number Prediction Head Number Cross-Attention F1inst. Accinst. F1img. Accimg. FPS 67.64 51.55 54.54 39.63 13.5 69.73 53.90 56.29 41.08 12.8 68.88 52.89 54.98 39.74 12.9 71.11 55.52 56.51 41.20 12.3
@@ 5.4 Limitations
Despite its strong performance on multi-target and no-target scenarios, NGDINO has several limitations. As illustrated by the qualitative examples in Figure 7, failure cases typically arise from challenging scenarios inherent to the RefDrone dataset. These failures can be categorized into three primary sources: (1) complexity in the referring expression that demands sophisticated contextual reasoning; (2) cluttered backgrounds that camouflage target objects; and (3) inherent difficulties in detecting very small-scale objects. Addressing these challenging, real-world conditions remains an important direction for future work.
@@ 6 Conclusion
In this work, we introduce RefDrone, a challenging benchmark for referring expression comprehension in drone scenes. The dataset is built using RDAnnotator, an efficient semi-automated annotation pipeline that combines LVLMs with human-in-the-loop verification to ensure high-quality annotations. Our extensive experiments reveal a substantial performance drop for existing state-of-the-art methods when evaluated on RefDrone, underscoring the benchmark’s difficulty and exposing key limitations in current REC approaches. Furthermore, we develop NGDINO to address the multi-target and no-target challenges in RefDrone. In the future, we aim to further enhance NGDINO to address additional challenges presented by RefDrone. We also plan to extend RefDrone to larger-scale scenarios and introduce additional tasks such as referring expression segmentation and tracking. We believe RefDrone will serve as a valuable benchmark for advancing research in drone-based REC tasks.
@@ Appendix A Appendix
  • Details of the baseline methods. (Appendix A.1)
@@
  • Details of our color categorization pipeline.
    (Appendix A.3)
  • Results with respect to different object scales.
    (Appendix A.4)
  • Ablation study on query length. (Appendix A.5)
  • Performance validation of the RDAnnotator framework. (Appendix A.6)
  • Results on standard RefCOCO/+/g datasets.
    (Appendix A.7)
  • Additional examples from the RefDrone dataset.
    (Appendix A.8)
  • Prompts and examples used in RDAnnotator.
    (Appendix A.9)
@@ A.2 Implementation details
NGDINO implementation. We train NGDINO using a two-stage procedure to ensure stability. Stage 1: Pre-training. We initialize the model with weights from a pre-trained GDINO [gdino], freezing all components except for the number prediction head for 5 epochs. This new head is then pre-trained on the RefDrone dataset. Stage 2: End-to-end Fine-tuning. After the head is pre-trained, we unfreeze the entire model and fine-tune it on our target dataset. This staged approach prevents the randomly initialized prediction head from destabilizing the well-trained detector backbone during the initial phases of training.
Zero-shot evaluation protocol. For all baseline models, we adhere to established evaluation practices to ensure fair comparisons. Specialized REC Methods: For GLIP [glip] and GDINO [gdino], we use the official model checkpoints and implementations provided within the MMDetection [mmdetection] framework. LVLMs: For a standardized and reproducible evaluation of LVLMs, we integrate our dataset into the VLMEvalKit framework [duan2024vlmevalkit]. Within this framework, we benchmark each model using its official, recommended prompt structure to ensure optimal performance.
Fine-tuning evaluation details. All fine-tuning experiments are conducted within the MMDetection [mmdetection] framework on 8 NVIDIA A100 GPUs. To ensure a fair comparison, we apply a consistent protocol across all models. We follow the original learning strategies and hyperparameter settings for each model with one critical modification: we disable random crop data augmentation. This is because random cropping can remove crucial spatial context or the target objects themselves in our position-sensitive referring expressions, introducing label noise and degrading performance.
@@
Color is a foundational attribute in the RefDrone dataset, present in 69% of all referring expressions. However, accurately identifying color is non-trivial due to challenges like illumination variance, occlusions, and semantic ambiguity (e.g., distinguishing "red" from "pink" or "orange"). To address this, we designed a hybrid color extraction pipeline that combines the efficiency of a specialized classifier with the reasoning capabilities of an LVLM. The pipeline consists of two stages:
(1) Classifier-based Proposal: A WideResNet-101 classifier [wideresnet] generates an initial color prediction. To create a high-quality training set for this classifier, we first generate labels programmatically using the HSV color space and then perform manual validation to correct noise and refine ambiguous cases.
(2) LVLM Verification: An LVLM verifier then assesses the classifier’s output. Using structured prompts, it reasons about the visual evidence to confirm the prediction or correct it, effectively resolving ambiguities caused by lighting or partial visibility.
The reliability of this hybrid approach enabled us to expand our vocabulary from an initial set of six primary colors (e.g., red, blue) to a more nuanced palette of twelve, including orange, pink, grey, and purple. The final distribution of these color terms is visualized in Figure 8.
@@
A.4 The results of different object scales.
To provide a granular analysis of model robustness to scale variation, a critical challenge in the RefDrone benchmark, we evaluated a representative set of high-performing methods (ACCinst. > 10%) on objects categorized as small, medium, and large. The results, presented in Table 7, reveal a stark performance gap between LVLMs and Specialized REC methods, particularly on small objects. Notably, even LVLMs like Qwen3-VL-235B struggle, achieving only 26.54% Accs. In contrast, among the fine-tuned specialized REC methods, our NGDINO-B achieves 44.08% Accs, 62.59% Accm, and 68.20% Accl.
Table 7: Performance comparison on small (Accs), medium (Accm), and large (Accl) objects. The models in the upper section are LVLMs evaluated in a zero-shot setting. The models in the lower section are specialized REC methods fine-tuned on the RefDrone training set.
Methods Params Accs Accm Accl Accinst. Rex-Omni [jiang2025detect] 3B 25.55 43.03 50.31 37.10 Qwen2.5-VL [bai2025qwen2] 3B 13.52 28.59 32.44 25.06 Qwen2.5-VL [bai2025qwen2] 7B 12.19 33.36 35.80 27.20 Qwen3-VL [qwen3technicalreport] 4B 18.53 42.66 54.23 32.31 Qwen3-VL [qwen3technicalreport] 8B 20.05 43.64 50.32 35.13 Qwen3-VL [qwen3technicalreport] 30B 24.21 50.48 54.68 39.98 Qwen3-VL [qwen3technicalreport] 235B 26.54 51.60 55.79 41.93 $\text{MiMo-VL}_{\text{RL}}$ [coreteam2025mimovltechnicalreport] 7B 2.41 13.69 15.65 11.22 GLM 4.1V [vteam2025glm45vglm41vthinkingversatilemultimodal] 9B 1.25 16.94 29.84 14.05 GLM 4.5V [vteam2025glm45vglm41vthinkingversatilemultimodal] 106B 7.70 27.83 29.17 21.97 $\text{DeepSeek-VL2}_{\text{Small}}$ 16B 5.83 27.64 30.36 21.34 DeepSeek-VL2 [wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels] 27B 2.66 18.58 18.57 14.88 DINO-XSeek [ren2024dino] 24.54 46.52 62.51 37.46 Seed1.5-VL [guo2025seed1] 11.11 38.39 50.83 27.84 Qwen3-VL-Plus [qwen3technicalreport] 25.34 49.96 56.18 41.32 GLIP-T [glip] 0.15B 21.08 48.82 54.35 40.39 GDINO-T [gdino] 0.17B 38.56 56.69 66.13 51.55 GDINO-B [gdino] 0.23B 40.01 59.96 67.94 53.95 NGDINO-T (Ours) 0.18B 42.48 60.42 67.90 55.52 NGDINO-B (Ours) 0.24B 44.08 62.59 68.20 57.22
A.5 Ablation study on query length.
Table 8 analyzes the impact of varying the query length. A minimal query length of 1 lacks the capacity to capture complex numerical information. Conversely, extending the query length to 100 increases parameter count and computational overhead, potentially leading to optimization challenges. Through these experiments, we determine that a query length of 10 provides an optimal trade-off.
Length F1inst. Accinst. F1img. Accimg. Params 1 70.20 54.44 55.82 40.56 1.58M 10 71.11 55.52 56.51 41.20 1.65M 100 70.44 54.72 55.95 40.73 2.34M
A.6 Performance of the RDAnnotator framework
To validate the effectiveness of our proposed annotation framework RDAnnotator, we evaluate RDAnnotator when repurposed as a complete, two-stage method for REC. To ensure a fair comparison, all methods operate on an identical set of initial object proposals generated by a first-stage Faster-RCNN detector [fasterrcnn] (18.0 mAP). The core of the evaluation lies in the second stage, where each method uses the referring expression to rank these proposals and identify the target. We benchmark RDAnnotator against two strong alternative second-stage approaches: (1) GPT-4o, which represents a powerful, single-step LVLM reasoning approach, and (2) ReCLIP [reclip], a representative CLIP-based ranker that relies on embedding similarity. As shown in Table 9, RDAnnotator substantially outperforms both, validating the efficacy of its structured, multi-step reasoning process. Results underscore RDAnnotator’s suitability for generating high-fidelity REC annotations.
Methods F1inst. Accinst. F1img. Accimg. ReCLIP [reclip] 24.62 14.04 11.58 6.15 GPT4-o 52.38 35.65 35.50 22.38 RDAnnotator 58.14 41.13 37.07 23.54
A.7 Results on RefCOCO/+/g datasets.
Since the RefCOCO, RefCOCO+, and RefCOCOg datasets contain only one instance per expression, the proposed NGDINO leverages the number branch primarily to address multi-instance and no-instance scenarios. As a result, the performance of NGDINO is relatively similar to that of GDINO on these datasets.
Table 10: Results on RefCOCO/+/g datasets.
RefCOCO RefCOCO+ RefCOCOg TestA TestB TestA TestB Val Test MDETR 90.4 82.67 85.52 72.96 83.35 83.31 GDINO-T 91.4 86.6 87.5 74.0 85.5 85.8 NGDINO-T 91.5 86.5 87.8 74.7 85.3 85.8
A.8 Dataset examples
@@
Figure 9: Dataset examples from RefDrone.
A.9 Prompts and examples for RDAnnotator